# LLLMs: A Data-Driven Survey of Evolving Research on Limitations of Large Language Models

### v1

@@ Abstract.
Large language model (LLM) research has grown rapidly, along with increasing concern about their limitations such as failures in reasoning, hallucinations, and limited multilingual capability. In this survey, we conduct a data-driven, semi-automated review of research on limitations of LLM (LLLMs) from 2022 to 2024 using a bottom-up approach. From a corpus of 250,000 ACL and arXiv papers, we identify 14,648 relevant papers using keyword filtering, LLM-based classification, validated against expert labels, and topic clustering (via two approaches, HDBSCAN+BERTopic and LlooM). We find that LLM-related research increases over fivefold in ACL and fourfold in arXiv. Since 2022, LLLMs research grows even faster, reaching over 30% of LLM papers by late 2024. Reasoning remains the most studied limitation, followed by generalization, hallucination, bias, and security. The distribution of topics in the ACL dataset stays relatively stable over time, while arXiv shifts toward safety and controllability (with topics like security risks, alignment, hallucinations, knowledge editing), and multimodality between 2022 and 2024. We release a dataset of annotated abstracts and a validated methodology, and offer a quantitative view of trends in LLM limitations research.
@@
ccs: Information systems Clustering and classificationccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Natural language processingccs: Computing methodologies Natural language generationccs: Computing methodologies Information extractionfootnotetext: © Aida Kostikova 2025. This is the author’s version of the work. It is posted here for your personal use. Not for redistribution. This is a pre-print of a paper submitted to ACM Computing Surveys.
@@ 1. Introduction
We observe four main results. i) LLLMs research has grown rapidly, outpacing even the growth of LLM research overall. The number of LLM-related papers has grown by a factor of 15 between 2022-2025, reaching almost 80% of all crawled ACL papers and roughly 30% of all crawled arXiv papers; LLLMs papers have increased by a factor of 28 in the same period, accounting for more than 30% of LLM papers in Q1 of 2025. ii) Within LLLMs research, reasoning limitations are the most prominent, with generalization, hallucination, bias, and security as further important concerns. iii) The distribution of limitations appears relatively stable in the ACL dataset, whereas the arXiv dataset shows a rise in concern for topics related to safety and controllability (e.g., Security Risks, Alignment Limitations, Knowledge Editing, Hallucination) as well as Multimodality. (iv) Despite substantial methodological differences between HDBSCAN and LlooM, we observe topical overlap in several of the biggest clusters (e.g., Reasoning, Hallucination, Security Risks) across both approaches, with broadly similar trend patterns, suggesting that these findings are reliable.
The contributions of this review to the field are i) a large-scale dataset of paper abstracts, tagged with limitation information, for further research, [^0] Data and code will be made publicly available. ii) an LLM-based paper annotation methodology, validated against human experts, iii) most importantly, quantitative insights into the evolution of LLLMs research covering the entire period 2022-2024 and early 2025, providing the first comprehensive overview of LLLMs research for this period.

### v2

@@ Abstract.
Large language model (LLM) research has grown rapidly, along with increasing concern about their limitations such as failures in reasoning, hallucinations, and limited multilingual capability. While prior reviews have addressed these issues, they often focus on individual limitations or consider them within the broader context of evaluating overall model performance. This survey addresses the gap by presenting a data-driven, semi-automated review of research on limitations of LLMs (LLLMs) from 2022 to 2025, using a bottom-up approach. From a corpus of 250,000 ACL and arXiv papers, we extract 14,648 relevant limitation papers using keyword filtering and LLM-based classification, validated against expert labels. Using topic clustering (via two approaches, HDBSCAN+BERTopic and LlooM), we identify between 7 and 15 prominent types of limitations discussed in recent LLM research across the ACL and arXiv datasets. We find that LLM-related research increases nearly sixfold in ACL and nearly fifteenfold in arXiv between 2022 and 2025, while LLLMs research grows even faster, by a factor of over 12 in ACL and nearly 28 in arXiv. Reasoning remains the most studied limitation, followed by generalization, hallucination, bias, and security. The distribution of topics in the ACL dataset stays relatively stable over time, while arXiv shifts toward safety and controllability (with topics like security risks, alignment, hallucinations, knowledge editing), and multimodality between 2022 and 2025. We offer a quantitative view of trends in LLM limitations research and release a dataset of annotated abstracts and a validated methodology, available at: github.com/a-kostikova/LLLMs-Survey.
@@
ccs: Information systems Clustering and classificationccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Natural language processingccs: Computing methodologies Natural language generationccs: Computing methodologies Information extraction
@@ 1. Introduction
We observe four main results. i) LLLMs research has grown rapidly, outpacing even the growth of LLM research overall. The number of LLM-related papers has grown by a factor of nearly 6 in ACL and nearly 15 in arXiv between 2022-2025, reaching almost 80% of all crawled ACL papers and roughly 30% of all crawled arXiv papers; LLLMs papers have increased even more sharply, by a factor of over 12 in ACL and 28 in arXiv, accounting for more than 30% of LLM papers in Q1 of 2025. ii) Within LLLMs research, reasoning limitations are the most prominent, with generalization, hallucination, bias, and security as further important concerns. iii) The distribution of limitations appears relatively stable in the ACL dataset, whereas the arXiv dataset shows a rise in concern for topics related to safety and controllability (e.g., Security Risks, Alignment Limitations, Knowledge Editing, Hallucination) as well as Multimodality. (iv) Despite substantial methodological differences between HDBSCAN and LlooM, we observe topical overlap in several of the biggest clusters (e.g., Reasoning, Hallucination, Security Risks) across both approaches, with broadly similar trend patterns, suggesting that these findings are reliable.
The contributions of this review to the field are i) a large-scale dataset of paper abstracts, tagged with limitation information, for further research,111github.com/a-kostikova/LLLMs-Survey ii) an LLM-based paper annotation methodology, validated against human experts, iii) most importantly, quantitative insights into the evolution of LLLMs research covering the entire period 2022-2024 and early 2025, providing the first comprehensive overview of LLLMs research for this period.